OpenStudy (anonymous):

Find the limit of (ln(x+1))/log2 x as x approaches to infinity.

OpenStudy (anonymous):

\[\lim_{x\to \infty}\frac{\ln(x+1)}{\ln(2x)}\]?

OpenStudy (primeralph):